Find the slope of the graph of the linear function of f
f(6) = -1, F(3)=8 I know the answer (-3) I just don't know how to get it.

Answer by Earlsdon(6294)

Try this:
Start with the slope-intercept form.
f(x) = mx+b Now write two equations using the given data.
f(6) = m(6)+b = -1
f(3) = m(3)+b = 8 Solve this system of equations by subtracting one from the other to find the slope, m.
6m + b = -1
-(3m + b = 8)
----------
3m = -9 Divide both sides by 3
m = -3
The slope is -3.